# Spatial Scale of Agglomeration and Dispersion: Number, Spacing, and the Spatial Extent of Cities

### v8

@@ E.1  Evaluating the impacts of local characteristics
Consider a spatial model with the indirect utility function $\bm{v}$. Let $a_{i}>0$ denote the innate characteristics of region $i$, and write $\bm{a}=(a_{i})_{i\in\mathcal{I}}$. For example, $a_{i}$ may represent the level of exogenous amenities or productivity in region ii [^25] All endogenous mechanisms related to agents’ spatial distribution $\bm{x}$, including endogenous amenities (e.g., Diamond-AER2016), are embedded in the indirect utility function $\bm{v}$ in our framework.. All regions are perfectly symmetric if $a_{i}=\bar{a}>0$ for all $i$. In this case, the uniform population distribution $\bar{\bm{x}}$ is an equilibrium.
@@ E.3  Numerical examples
This section provides numerical examples for Proposition˜5. To introduce exogenous asymmetry, we multiply the utility in region $1$ by $a_{1}\geq 1$, whereas we let $a_{i}=1$ for all $i\neq 1$. We consider the Krugman model and Allen–Arkolakis model, and basic model parameters are set to be the same as Fig.˜C.3 and Fig.˜C.3 except that region $1$ has an exogenous advantage.
Figure˜2A reports equilibrium paths of $x_{1}$ for the Allen–Arkolakis model (Type L) under the uniqueness of the equilibrium. The curves depict region $1$’s population share, $x_{1}$, at stable equilibria against $\phi$. Four incremental settings $a_{1}\in\{1.000,1.001,1.005,1.010\}$ are considered, including the baseline case with no location-fixed advantage ($a_{1}=1.000$). We have $\delta^{\prime}(\Theta)>0$ for all $\Theta\in(0,1)$ and see that $x_{1}-\bar{x}>0$ when $a_{1}>1$ and $x_{1}-\bar{x}$ increases as $a_{1}$ increases, which are intuitive. Additionally, $x_{1}-\bar{x}$ decreases as $\phi$ increases. We confirm that $\rho(\phi)>0$ and $\rho^{\prime}(\phi)<0$ for all $\phi$.
@@ C Type G
Figure E.2: Population share of the advantageous region $1$ and covariance $\rho$
@@ E.4  Proof of Proposition˜5
where we make the dependence of $\bm{v}$ on $\bm{a}$ explicit, $\bar{v}(\bm{x},\bm{a})\equiv\sum_{i\in\mathcal{I}}v_{i}(\bm{x},\bm{a})x_{i}$ is the average utility, and $\bm{1}$ is $N$-dimensional all-one vector. The pair $(\bar{\bm{x}},\bar{\bm{a}})$ is a solution to Eq.˜E.7. Suppose that there is a spatial equilibrium nearby $\bar{\bm{x}}$ when $\bm{a}$ is marginally different from $\bar{\bm{a}}$. Let $\bm{x}(\bm{a})$ denote the perturbed version of $\bar{\bm{x}}$, which is a function in $\bm{a}$. We assume $\bar{\bm{x}}$ is stable so that studying a perturbed version of it makes sense.
@@ F.2.1  Krugman-JPE1991 model
There are two types of workers, mobile and immobile, and their total masses are $1$ and LL, respectively. $\bm{x}\equiv(x_{i})_{i\in\mathcal{I}}$ is the distribution of mobile workers. Each worker supplies one unit of labor inelastically.
@@
that is, $u_{i}=(C_{i}^{\mathrm{M}})^{\mu}(C_{i}^{\mathrm{A}})^{1-\mu}$ where $\mu\in(0,1)$ is the constant expenditure of the latter. With free trade in the A-sector, the wage of the immobile worker is equalized, and we normalize it to unity by taking A-sector goods as the numéraire. Consequently, region $i$’s expenditure on the M-sector goods is given by $e_{i}=\mu(w_{i}x_{i}+l_{i})$ where $l_{i}$ denotes the mass of immobile workers in region $i$.
In the M-sector, to produce $q$ units, a firm requires $\alpha+\beta q$ units of mobile labor. Profit maximization of firms yields the price of differentiated goods produced in region $i$ and exported to $j$ as $p_{ij}=\frac{\sigma\beta}{\sigma-1}w_{i}\tau_{ij}$, which in turn determines gravity trade flow from $j$ to $i$. That is, when $X_{ij}$ denotes the price of M-sector goods produced in region $i$ and sold in region $j$, $X_{ij}=m_{ij}e_{j}$ where the share $m_{ij}\in(0,1)$ is defined by Eq.˜F.6 with $\phi_{ij}\equiv\tau_{ij}^{1-\sigma}$. The proximity matrix is thus $\phi_{ij}]=[\tau_{ij}^{1-\sigma}]$.
@@ F.2.2  Helpman-Book1998 and Redding-Sturm-AER2008 model
There are two variants for assumptions on how housing stocks are owned: public landownership (PL) and local landownership (LL). Helpman-Book1998 supposes PL in which housing stocks are equally owned by all workers; the income of a worker in region $i$ is the sum of the wage and an equal dividend $r>0$ of the total rental revenue in the economy. However, ottaviano2002agglomeration, Murata-Thisse-JUE2005, and Redding-Sturm-AER2008 assumed that housing stocks are locally owned (i.e., LL). The income of a worker in region $i$ is the sum of the wage and an equal dividend of rental revenue in each region. In fact, the model by Redding-Sturm-AER2008 is the LL version of the Helpman model.
@@
Also, the market wage is given as the solution for Eq.˜F.5. For the PL case, we set $r=1$ for normalization. For the LL case, $\bm{w}(\bm{x})$ is uniquely given up to normalization. The indirect utility function is, with $\Delta_{i}\equiv\sum_{j\in\mathcal{I}}x_{j}w_{j}^{1-\sigma}\phi_{ji}$ and $r>0$,
@@
For the LL case, $\mathbf{W}_{x}$ is given in Example˜6 and we obtain
@@
Equilibrium is unique when $\gamma\sigma=(1-\mu)\sigma>1$ (Redding-Sturm-AER2008). For both PL and LL, this condition implies that $\Omega^{\sharp}(\Theta)<0$ for all $\Theta\in(0,1)$. ∎
@@
The regional model in §3 of Redding-Rossi-Hansberg-ARE2017 is a variant of the Helpman model with LL, in which variable input of mobile labor depends on region $i$ (i.e., productivity differs across regions). The cost function of firms in region $i$ is $C_{i}(q)=w_{i}(\alpha+\beta_{i}q)$. The market equilibrium condition for this case is, with $a_{i}\equiv\beta_{i}^{1-\sigma}>0$, given by
@@
where $\Omega^{\sharp}(\Theta)$ is that for the LL case Eq.˜F.35. If $(1-\mu)\sigma>1$, $\delta^{\prime}(\Theta)>0$ for all $\Theta$. ∎
@@ F.2.3  Pfluger-Suedekum-JUE2008 model
where $\Delta_{i}=\sum_{j\in\mathcal{I}}\phi_{ji}x_{j}$, and $l_{i}$ and $a_{i}$ denote the mass of immobile workers and amount of housing stock in region $i$, respectively. The nominal wage in region $i$ is given by Eq.˜F.39. Let $l_{i}=l$ and $a_{i}=a$ for all $i$. Then, we see that $\mathbf{V}=\frac{1}{\bar{v}}\Omega^{\sharp}(\mathbf{D})$ with

### v9

@@ E.1  Evaluating the impacts of local characteristics
Consider a spatial model with the indirect utility function $\bm{v}$. Let $a_{i}>0$ denote the innate characteristics of region $i$, and write $\bm{a}=(a_{i})_{i\in\mathcal{I}}$. For example, $a_{i}$ may represent the level of exogenous amenities or productivity in region ii [^25] All endogenous mechanisms related to agents’ spatial distribution $\bm{x}$, including endogenous amenities (e.g., Diamond-AER2016), are embedded in the indirect utility function $\bm{v}$ in our framework.. All regions are perfectly symmetric if $a_{i}=\bar{a}>0$ for all ii. In this case, the uniform population distribution $\bar{\bm{x}}$ is an equilibrium.
@@ E.3  Numerical examples
This section provides numerical examples for Proposition˜5. To introduce exogenous asymmetry, we multiply the utility in region 11 by $a_{1}\geq 1$, whereas we let $a_{i}=1$ for all $i\neq 1$. We consider the Krugman model and Allen–Arkolakis model, and basic model parameters are set to be the same as Fig.˜C.3 and Fig.˜C.3 except that region 11 has an exogenous advantage.
Figure˜2A reports equilibrium paths of $x_{1}$ for the Allen–Arkolakis model (Type L) under the uniqueness of the equilibrium. The curves depict region 11’s population share, $x_{1}$, at stable equilibria against $\phi$. Four incremental settings $a_{1}\in\{1.000,1.001,1.005,1.010\}$ are considered, including the baseline case with no location-fixed advantage ($a_{1}=1.000$). We have $\delta^{\prime}(\Theta)>0$ for all $\Theta\in(0,1)$ and see that $x_{1}-\bar{x}>0$ when $a_{1}>1$ and $x_{1}-\bar{x}$ increases as $a_{1}$ increases, which are intuitive. Additionally, $x_{1}-\bar{x}$ decreases as $\phi$ increases. We confirm that $\rho(\phi)>0$ and $\rho^{\prime}(\phi)<0$ for all $\phi$.
@@ C Type G
Figure E.2: Population share of the advantageous region 11 and covariance $\rho$
@@ E.4  Proof of Proposition˜5
where we make the dependence of $\bm{v}$ on $\bm{a}$ explicit, $\bar{v}(\bm{x},\bm{a})\equiv\sum_{i\in\mathcal{I}}v_{i}(\bm{x},\bm{a})x_{i}$ is the average utility, and $\bm{1}$ is NN-dimensional all-one vector. The pair $(\bar{\bm{x}},\bar{\bm{a}})$ is a solution to Eq.˜E.7. Suppose that there is a spatial equilibrium nearby $\bar{\bm{x}}$ when $\bm{a}$ is marginally different from $\bar{\bm{a}}$. Let $\bm{x}(\bm{a})$ denote the perturbed version of $\bar{\bm{x}}$, which is a function in $\bm{a}$. We assume $\bar{\bm{x}}$ is stable so that studying a perturbed version of it makes sense.
@@ F.2.1  Krugman-JPE1991 model
There are two types of workers, mobile and immobile, and their total masses are 11 and $L$, respectively. $\bm{x}\equiv(x_{i})_{i\in\mathcal{I}}$ is the distribution of mobile workers. Each worker supplies one unit of labor inelastically.
@@
that is, $u_{i}=(C_{i}^{\mathrm{M}})^{\mu}(C_{i}^{\mathrm{A}})^{1-\mu}$ where $\mu\in(0,1)$ is the constant expenditure of the latter. With free trade in the A-sector, the wage of the immobile worker is equalized, and we normalize it to unity by taking A-sector goods as the numéraire. Consequently, region $i$’s expenditure on the M-sector goods is given by $e_{i}=\mu(w_{i}x_{i}+l_{i})$ where $l_{i}$ denotes the mass of immobile workers in region ii.
In the M-sector, to produce $q$ units, a firm requires $\alpha+\beta q$ units of mobile labor. Profit maximization of firms yields the price of differentiated goods produced in region $i$ and exported to $j$ as $p_{ij}=\frac{\sigma\beta}{\sigma-1}w_{i}\tau_{ij}$, which in turn determines gravity trade flow from $j$ to ii. That is, when $X_{ij}$ denotes the price of M-sector goods produced in region $i$ and sold in region $j$, $X_{ij}=m_{ij}e_{j}$ where the share $m_{ij}\in(0,1)$ is defined by Eq.˜F.6 with $\phi_{ij}\equiv\tau_{ij}^{1-\sigma}$. The proximity matrix is thus $\phi_{ij}]=[\tau_{ij}^{1-\sigma}]$.
@@ F.2.2  Helpman-Book1998 and Redding-Sturm-AER2008 model
There are two variants for assumptions on how housing stocks are owned: public landownership (PL) and local landownership ($L$). Helpman-Book1998 supposes PL in which housing stocks are equally owned by all workers; the income of a worker in region $i$ is the sum of the wage and an equal dividend $r>0$ of the total rental revenue in the economy. However, ottaviano2002agglomeration, Murata-Thisse-JUE2005, and Redding-Sturm-AER2008 assumed that housing stocks are locally owned (i.e., $L$). The income of a worker in region $i$ is the sum of the wage and an equal dividend of rental revenue in each region. In fact, the model by Redding-Sturm-AER2008 is the $L$ version of the Helpman model.
@@
Also, the market wage is given as the solution for Eq.˜F.5. For the PL case, we set $r=1$ for normalization. For the $L$ case, $\bm{w}(\bm{x})$ is uniquely given up to normalization. The indirect utility function is, with $\Delta_{i}\equiv\sum_{j\in\mathcal{I}}x_{j}w_{j}^{1-\sigma}\phi_{ji}$ and $r>0$,
@@
For the $L$ case, $\mathbf{W}_{x}$ is given in Example˜6 and we obtain
@@
Equilibrium is unique when $\gamma\sigma=(1-\mu)\sigma>1$ (Redding-Sturm-AER2008). For both PL and $L$, this condition implies that $\Omega^{\sharp}(\Theta)<0$ for all $\Theta\in(0,1)$. ∎
@@
The regional model in §3 of Redding-Rossi-Hansberg-ARE2017 is a variant of the Helpman model with $L$, in which variable input of mobile labor depends on region $i$ (i.e., productivity differs across regions). The cost function of firms in region $i$ is $C_{i}(q)=w_{i}(\alpha+\beta_{i}q)$. The market equilibrium condition for this case is, with $a_{i}\equiv\beta_{i}^{1-\sigma}>0$, given by
@@
where $\Omega^{\sharp}(\Theta)$ is that for the $L$ case Eq.˜F.35. If $(1-\mu)\sigma>1$, $\delta^{\prime}(\Theta)>0$ for all $\Theta$. ∎
@@ F.2.3  Pfluger-Suedekum-JUE2008 model
where $\Delta_{i}=\sum_{j\in\mathcal{I}}\phi_{ji}x_{j}$, and $l_{i}$ and $a_{i}$ denote the mass of immobile workers and amount of housing stock in region $i$, respectively. The nominal wage in region $i$ is given by Eq.˜F.39. Let $l_{i}=l$ and $a_{i}=a$ for all ii. Then, we see that $\mathbf{V}=\frac{1}{\bar{v}}\Omega^{\sharp}(\mathbf{D})$ with